# Early Stages of Self-Healing at Tungsten Grain Boundaries from Ab Initio Machine Learning Simulations

### v1

@@ 2.3 Dynamics and mechanism of defect migration to the GB
After comparing structural properties and defect energetics, we now turn to study defect evolution in a W110delimited-⟨⟩110$\langle$ 110$\rangle$ 110 ⟩/W112delimited-⟨⟩112$\langle$ 112$\rangle$ 112 ⟩ GB supercell containing 54,000 atoms. This size was chosen as a compromise between computational cost and the ability to reproduce bulk-like conditions within the grains. To study this process, we introduced 500, 750, and 1,000 defects (Frenkel pairs) into the system to simulate typically expected post-cascade defect densities (hereafter referred to as the cases of low, medium and high density of introduced defects, respectively), followed by thermalization at the appropriate temperatures (300 K, 650 K, and 1000 K) for 8 ps, a value that falls within the range commonly employed for this type of simulation. Fig. 2 shows the initial (0 ps) and final (100 ps) configurations for the third case, i.e., high density of defects, obtained using MLIP-W (Fig. 2 A), the Tersoff potential (Fig. 2 B), and the EAM potential (Fig. 2 C), at 300 K (top panels) and 1000 K (bottom panels). Atoms in Fig. 2 are color-coded according to their potential energy difference ($\Delta E_{\text{pot}}$) relative to the bulk for each potential. We only show atoms with energies above 0.5 eV for 300 K and 0.75 eV for 1000 K, in order to highlight defected or strained regions. At high temperatures, the increase in the minimum energy value accounts for thermal vibrations. In addition to the atomic configurations, histograms in Fig. 2 illustrate the distribution of potential energies. In these histograms, the red curve corresponds to the initial distribution (0 ps), while the blue curve represents the final distribution (100 ps). Moreover, atoms not visualized in the atomic configurations correspond to those within the gray shaded area of the histogram.
@@
At 1000 K, the MLIP-W (Fig. 2 A) maintains a high level of structural order at 0 ps, with atoms having $\Delta E_{\text{pot}}>0.75$ eV located in definite defect-rich regions, mainly forming dumbbell configurations. The percentage of these atoms is 3.0%, slightly lower than at 300 K. The number of atoms with $\Delta E_{\text{pot}}>2$ eV increases, probably due to the enhancement of thermal vibrations. As shown in the lower part of Fig. 2, after 100 ps, the migration of SIAs to the GB becomes more efficient and the number of atoms mainly located at the GB with $\Delta E_{\text{pot}}>0.75$ eV decreases to 1.4%. This defect annihilation, reflected in the histogram, demonstrates the ability of the potential to capture defect dynamics at elevated temperatures and is consistent with experimental observations of denuded zones near GBs [32]. Similarly to the situation at 300 K, calculations with the Tersoff potential (Fig. 2 B) show that the system remains disordered at 1000 K with no clear defect structures such as dumbbells. The remaining atoms do not correspond to defected regions, indicating that the Tersoff potential struggles to distinguish defects from thermal oscillations, leading to a less reliable defect evolution representation. In the final state, only 1.8% of atoms have $\Delta E_{\text{pot}}>0.75$ eV, most of them at the GB, though this is misleading as the potential fails to associate high-energy atoms to defects. Calculations with the EAM potential (Fig. 2 C) indicate that high-energy atoms remain localized in defect regions, with 2.6% of atoms exhibiting $\Delta E_{\text{pot}}>0.75$ eV. This number of atoms is nearly double that calculated with the MLIP-W. The EAM potential also retains some crowdion-like structures very close to the GB, which is unexpected given their proximity. The inability of the Tersoff and EAM potentials to accurately capture defect migration and stabilization near GBs reveals significant limitations that could lead to misinterpretation in radiation-induced damage studies, especially when accurate modeling of defect sinks is critical to predict material performance.
The results reported above are supported by direct visualizations from the MD movies provided in Supplementary Note 2, which show the dynamics captured in 40 snapshots over 100 ps for each IP (see Figs. S2-S10). The time-evolution snapshots shown in Fig. 3 A and B illustrate the dynamic migration of $\langle 111\rangle$ dumbbells from both the $\langle 112\rangle$ (Fig. 3 A) and $\langle 110\rangle$ (Fig. 3 B) grain interiors to the GB at 300 K, respectively. These snapshots show that the stable dumbbell configurations facilitate an efficient defect transport mechanism in which SIAs undergo sequential displacements along preferential crystallographic directions, effectively propagating defects toward the GB, as previously mentioned. The top and bottom rows, in the same figure, show cross sections along $\overline{1}$} (XZ plane) and $\bar{1}$0} (YZ plane), providing complementary perspectives on this migration mechanism. The crystallographic directions shown on the right side of Fig. 3 allow direct correlation with the underlying lattice. This efficient transport process, in which defects are directed to the GB as an effective sink, highlights the ability of the MLIP-W to accurately capture the mechanism of defect evolution, in agreement with DFT and Monte Carlo predictions [61].
@@
In addition, for a more objective analysis, we define a region around the GB whose z boundaries do not intersect any atomic layers, thus preventing thermal fluctuations from being mistaken for actual migrations. By comparing the atoms initially in this region with those at the end of the simulation, we identify new atoms (tracked via their unique IDs) as those that have migrated to the GB. The exact values are shown in Table S1 of Supplementary Note 3. At low temperatures (e.g., 300 K), differences among the three IPs become particularly pronounced: the number of SIAs migrating to the GB with the MLIP-W is around ten times higher than that calculated with the Tersoff and EAM IPs. These results agree much better with previously reported DFT data, which have shown that the migration barriers for SIAs in tungsten are close to zero [62], and with Monte Carlo simulations predicting that this migration can occur at temperatures as low as 5-10 K [61]. They are also in agreement with experimental data reported for nanostructured W samples sequentially implanted with C, which show that these materials exhibit greater H retention than their bulk counterpart, as the SIAs migrate to the GB, leaving vacancies in the grain interior where H is accumulated [63, 64]. At intermediate and high temperatures, the differences between the potentials become more subtle. At 650 K, the number of defects that have migrated to the GB calculated with MLIP-W and Tersoff potentials are very similar on average, while that calculated with the EAM potential is a factor of two lower. At 1000 K, the number of defects that have migrated to the GB calculated with the MLIP-W is about 2.5 times higher than that calculated with the Tersoff or the EAM potentials. These results prove the high fidelity of the MLIP-W in capturing both the energetics and kinetics of SIAs migrating to GB, as well as the limitations of the Tersoff and EAM IPs in describing these processes.
@@ 2.4 GB structure after defect migration
Having determined the number of defects migrating to the GB, we now delve into the atomistic structure of the interface given by each IP. Fig. 4 shows the potential energies of the GB atoms relative to the bulk as determined with each potential. At 0 K, for pristine samples (top row), all three potentials converge to a similar interface structure with a periodic pentagonal motif (yellow) and a comparable equilibrium geometry. However, the maximum GB-bulk energy difference depends on the selected IPs being 0.77 eV (for the MLIP-W), 0.91 eV (for the EAM potential), and 1.05 eV (for the Tersoff potential). Based on previous DFT studies [65, 66], the interface naturally forms GB-aligned grooves, which serve as defect sinks. At temperatures from 300 K to 1000 K, the GB structure estimated with the MLIP-W remains almost unchanged, while the motifs begin to blur when the other IPs are used (not shown in Fig. 4). When we further analyze the GB after introducing Frenkel pairs (high density case) and allow the system to evolve at different temperatures, diverse changes in defect behavior were observed. The middle (300 K) and bottom (1000 K) rows of Fig. 4 show the final GB configurations under defect implantation conditions comparable to those in Fig. 2. The DFT findings mentioned above indicate that SIAs arrange in rows along these grooves because these regions have a lower defect formation energy than the bulk, largely due to the larger space available for accommodation. For the MLIP-W, SIAs are observed to effectively reach the GB and align along these grooves. In particular, at finite temperatures, atoms near both the W112delimited-⟨⟩112$\langle$ 112$\rangle$ 112 ⟩ and W110delimited-⟨⟩110$\langle$ 110$\rangle$ 110 ⟩ layers adjacent to the SIAs row shift into slightly arc-shaped crowdions to minimize the energy of the system. These regions are marked as green hexagons in Fig. 4 to illustrate the characteristic atomic packing. This phenomenon is explained in more detail in Supplementary Note 4, where Fig. S11 clearly illustrates the formation of the crowdion. At low temperatures, such distribution is also captured, up to a certain degree, by the Tersoff and the EAM IPs. However, the ability of these empirical IPs to reproduce the GB structure decreases with increasing temperature to such an extent that they predict a compromised integrity of the GB, which is unexpected at the temperatures studied (up to 1000 K).
@@ 2.5 Descriptor-based analysis of defects
To further validate and analyze the performance of the MLIP-W in describing defect evolution at W110delimited-⟨⟩110$\langle$ 110$\rangle$ 110 ⟩/W112delimited-⟨⟩112$\langle$ 112$\rangle$ 112 ⟩ GBs, we used the Fingerprinting and Visualization Analyzer of Defects (FaVAD) [39]. This approach is based on DVs and allows us to characterize defects based on their local atomic environments. Although the DeePMD-kit [42, 43, 67] generates its own DVs, their high dimensionality made them impractical for our 54,000-atom cell, and we instead use the multi-body descriptor framework called ‘Smooth Overlap of Atomic Positions’ (SOAP) [40] to compute a reduced set of descriptors (see Section 4.4). In this fingerprinting method, we measure the Euclidean distance, denoted as $d^{M}(T)$, between the DV of each atom in the defect sample, $\vec{q}^{\,i}$, and the mean DV of a defect-free (bulk) reference configuration at a certain temperature $T$, $\vec{v}(T)$. This distance difference between two local atomic environments forms the basis for a probabilistic interpretation of the results, expressed as $p\bigl{(}\vec{q}^{\,i}\mid\vec{v}(T)\bigr{)}$. We find that the distribution of these normalized distances is well approximated by a chi-squared distribution, $\chi^{2}_{K}$, with $K$ degrees of freedom. Mathematically, this is expressed as
@@
where $\Gamma$ is the Gamma function. In our analysis, as observed in Fig. S13, $K=4$ provides an optimal fit to the distributions [39]. This probabilistic framework allows defining a temperature-dependent threshold to differentiate genuine defects from thermal oscillations. To establish this threshold, first the bulk reference DVs are computed at each simulation temperature. Since defects in BCC materials predominantly form $\langle 111\rangle$ dumbbell configurations [68], reference DVs for thermally equilibrated interstitials are also determined. Note that reference DVs are determined individually for each potential. By mapping distances relative to the bulk and interstitial reference DVs for a pure bulk configuration at a given temperature, a distribution that defines the bulk region in 2D space is obtained. Thus, when analyzing configurations containing defects or GBs, points outside this region are considered defects. This 2D methodology represents an advance over that used in Ref. [39], as it allows for a more accurate identification of crowdion-like interstitials, which are the most likely structures. Further details of this methodology, including how the area thresholds are determined, are given in Supplementary Note 5; in particular, Fig. S12 shows the distributions found for a 54,000-atom bulk system thermalized at each temperature and for each potential. The DV distributions reveal that while the MLIP-W broadens uniformly at higher temperatures, empirical IPs—particularly the Tersoff IP—develop bifurcated structures at 1000 K that are indicative of distinct structural motifs, a behavior likely attributable to an artifact of the potential. Exact thresholds can be found in Table S2 of the Supplementary Note 6.
@@
Fig. 5 shows DV distributions (left) and simulation cells (right) for the high density of introduced defects case, both at 300 K and 1000 K. Atoms are color coded by environment: SIAs (light purple), vacancies (orange), and unidentified defects (dark purple), with bulk atoms omitted for the sake of clarity. Two new "sheets" emerge: the lower one for SIAs and the upper one for the GB and unidentified defects. In Fig. S13 of Supplementary Note 7, a detailed comparison of the DV distributions for bulk, pristine GB, and defected GB, each thermalized at 300 K, is presented. At 300 K, all potentials exhibit similar distributions, though the EAM IP shows a slight “V” shape. The MLIP-W produces fewer SIAs and defects, resulting in fewer points in the 3D cell. The EAM potential behaves similarly, while the Tersoff potential, again, struggles to distinguish defects. At 1000 K, differences become more pronounced. The distribution found with the MLIP-W expands like in the bulk case, sharply reducing interstitials (by a factor of ten) and unidentified defects (by a factor of three) due to a broader bulk region filtering thermally vibrating atoms. The Tersoff IP shows an even steeper decline, from 4,129 defects at 300 K to only 185, as defects become indistinguishable from thermal noise. EAM behaves similarly to the MLIP-W but introduces a third “sheet”, suggesting additional defective regions. It also retains twice as many SIAs as the MLIP-W (253 vs. 138), likely due to fewer migration events.
@@ 2.6 Quantifying the temperature dependence of defect migration
Fig. 6 shows the average number of defects calculated with each IP as a function of temperature and initial defect density. For these counts, the GB is considered as a defect sink and therefore the atoms located at the GB are excluded. Full numerical values are provided in Supplementary Note 8, where Tables S3, S4, and S5 list the defect counts for the MLIP-W, the Tersoff, and the EAM potentials, respectively. At 300 K, in agreement with our previous observations, the MLIP-W consistently yields the lowest number of defects, whereas the Tersoff and the EAM potentials predict higher counts. Specifically, for the low density of introduced defects case, the number of defects calculated with the Tersoff and the EAM potentials is 1.7 and 1.2 times larger than that calculated with the MLIP-W, respectively. For the medium density of introduced defects case, the number of defects calculated with the Tersoff and the EAM IPs is 1.7 and 2.3 times higher than that calculated with the MLIP-W. Finally, for the high density of defects introduced case, the number of defects calculated with the Tersoff and the EAM IPs is 1.8 and 1.4 times larger than that calculated with the MLIP-W. These trends indicate that, at low temperatures, both the Tersoff and the EAM IPs yield more residual defects in the grain interior than the MLIP-W, evidencing the superior ability of the MLIP-W to capture defect annihilation and/or reorganization at the GB. As temperature increases, the differences between the IPs become less pronounced. The EAM potential appears to follow a trend similar to the MLIP-W, albeit with consistently higher defect counts, whereas the slope of the Tersoff curves is considerably steeper, likely reflecting its inability to accurately capture defect behavior at high temperatures, resulting in defect counts even lower than those predicted by the MLIP-W.
@@
A natural way to validate the temperature dependence of these defect count curves is to interpret them in the light of the kinetic rate theory by Arrhenius [69, 60]. Assuming that SIAs, once formed, diffuse to the GBs, that act as sinks for them, the annihilation rate or net decrease of SIAs at the GB with regard to the grain interior can be considered an effective measure of their migration. Such annihilation rate $v$ follows a classical Arrhenius form,
@@
where $\Delta n_{m}$ is the variation in the number of defects inside the grain (excluding vacancies) in a time $\Delta t$, $N$ is the total number of atoms, $S_{m}$ is the migration entropy, $E_{m}$ is the migration energy, and $k_{B}$ is the Boltzmann constant. The temperature-independent prefactor $\exp\bigl{(}S_{m}/k_{B}\bigr{)}$ represents the entropic contribution to the formation of an activated state, while $\exp\bigl{(}-E_{m}/k_{B}T\bigr{)}$ captures the energetic barrier to defect migration. Considering that $\Delta t$ is constant and equal to 100 ps in all simulations, we can fit Eq. (2) to our $\Delta n_{m}/N$ vs. $T$ data to extract $E_{m}$. We find that the MLIP-W yields a migration energy of $0.048\pm 0.001$ eV, which is in good agreement with the typical $\sim 0.05$ eV range reported for SIAs in W—for instance, 0.054 eV [60] or 0.046 eV [70]. For the EAM potential, the fits lead to a slightly lower migration energy value of $0.043\pm 0.001$ eV. The migration energy as fitted with the Tersoff potential has a much larger value than that calculated by the other potentials ($0.082\pm 0.008$ eV). These results are additional evidence of the inaccuracy of the Tersoff potential to distinguish defects from thermal noise. The good agreement of $E_{m}$ calculated with the MLIP-W and the reference data highlights its to capture defect kinetics and GB-mediated self-healing mechanisms in W under conditions where previously reported empirical IPs do not perform well.

### v2

@@ 2.3 Dynamics and mechanism of defect migration to the GB
After comparing structural properties and defect energetics, we now turn to study defect evolution in a $\langle 110\rangle/W$$\langle 112\rangle$ GB supercell containing 54,000 atoms. This size was chosen as a compromise between computational cost and the ability to reproduce bulk-like conditions within the grains. For the large-scale MD simulations that follow, we restrict the comparison to the two widely used empirical potentials, EAM and Tersoff, which are far less computationally demanding and, to date, have been used more broadly than the more recent MLIPs.
To study this process, we introduced 500, 750, and 1,000 defects (Frenkel pairs) into the system to simulate typically expected post-cascade defect densities (hereafter referred to as the cases of low, medium, and high density of introduced defects, respectively), followed by thermalization at the appropriate temperatures (300 K, 650 K, and 1000 K) for 8 ps, a value that falls within the range commonly employed for this type of simulation. Fig. 3 shows the initial (0 ps) and final (100 ps) configurations for the third case, i.e., high density of defects, obtained using MLIP-W (Fig. 3 A), the Tersoff potential (Fig. 3 B), and the EAM potential (Fig. 3 C), at 300 K (top panels) and at 1000 K (bottom panels). Atoms in Fig. 3 are color-coded according to their potential energy difference ($\Delta E_{\text{pot}}$) relative to the bulk for each potential. We only show atoms with energies above 0.5 eV at 300 K and 0.75 eV at 1000 K, in order to highlight defected or strained regions. At high temperatures, the increase in the minimum energy value accounts for thermal vibrations. In addition to the atomic configurations, histograms in Fig. 3 illustrate the distribution of potential energies. In these histograms, the red curve corresponds to the initial distribution (0 ps), while the blue curve represents the final distribution (100 ps). Moreover, atoms not visualized in the atomic configurations correspond to those within the gray shaded area of the histogram.
@@
At 1000 K, the MLIP-W (Fig. 3 A) maintains a high level of structural order at 0 ps, with atoms having $\Delta E_{\text{pot}}>0.75$ eV located in definite defect-rich regions, mainly forming dumbbell configurations. The percentage of these atoms is 3.0%, slightly lower than at 300 K. The number of atoms with $\Delta E_{\text{pot}}>2$ eV increases, probably due to the enhancement of thermal vibrations. As shown in the lower part of Fig. 3, after 100 ps, the migration of SIAs to the GB becomes more efficient and the number of atoms mainly located at the GB with $\Delta E_{\text{pot}}>0.75$ eV decreases to 1.4%. This defect annihilation, reflected in the histogram, demonstrates the ability of the potential to capture defect dynamics at elevated temperatures and is consistent with experimental observations of denuded zones near GBs [32]. Similarly to the situation at 300 K, calculations with the Tersoff potential (Fig. 3 B) show that the system remains disordered at 1000 K with no clear defect structures such as dumbbells. The remaining atoms do not correspond to defected regions, indicating that the Tersoff potential leads to atomic configurations in which defect-related and thermal displacement environments overlap significantly, which complicates defect identification. In the final state, only 1.8% of the atoms exhibit $\Delta E_{\text{pot}}>0.75$ eV, most of which are located at the GB. However, this result should be viewed with caution since, as mentioned just above, high potential energy does not reliably indicate an actual defect. Finally, calculations with the EAM potential (Fig. 3 C) indicate that high-energy atoms remain localized in defect regions, with 2.6% of atoms exhibiting $\Delta E_{\text{pot}}>0.75$ eV. This number of atoms is nearly double that calculated with the MLIP-W. The EAM potential also retains some crowdion-like structures very close to the GB, which is unexpected given their proximity. The superior performance of MLIP-W likely originates from its extensive DFT-based training dataset. This enables MLIP-W to accurately describe stable defect configurations, such as $\langle 111\rangle$ interstitial dumbbells, as well as subtle transition pathways and energy barriers between intermediate states. By contrast, empirical potentials typically lack explicit fitting to these non-obvious energetic features, resulting in a less accurate, rougher PES. Such discrepancies can lead to misinterpretations in radiation-induced damage studies, especially when precise modeling of defect sinks is necessary to predict material performance.
[…]
@@
In addition, for a more objective analysis, we define a region around the GB whose z boundaries do not intersect any atomic layers, thus preventing thermal fluctuations from being mistaken for actual migrations. By comparing the atoms initially in this region with those at the end of the simulation, we identify new atoms (tracked via their unique IDs) as those that have migrated to the GB. The exact values are shown in Table S1 of Supplementary Note 3. At low temperatures (e.g., 300 K), differences among the three IPs become particularly pronounced: the number of SIAs that migrate to the GB, as calculated using the MLIP-W, is approximately ten times higher than the number calculated using the Tersoff and EAM IPs. The results provided by the MLIP-W agree much better with previously reported DFT data, which have shown that the migration barriers for SIAs in tungsten are close to zero [64], and with Monte Carlo simulations predicting that this migration can occur at temperatures as low as 5-10 K [63]. They are also in agreement with experimental data reported for nanostructured W samples sequentially implanted with C, which show that these materials exhibit greater H retention than their bulk counterpart, as the SIAs migrate to the GB, leaving vacancies in the grain interior where H is accumulated [65, 66]. At intermediate and high temperatures, the differences between the potentials become subtler. At 650 K, the number of defects that have migrated to the GB calculated with MLIP-W and Tersoff potentials is very similar on average, while that calculated with the EAM potential is a factor of two lower. At 1000 K, the number of defects that have migrated to the GB calculated with the MLIP-W is about 2.5 times higher than that calculated with the Tersoff or the EAM potentials. These results prove the high fidelity of the MLIP-W in capturing both the energetics and kinetics of SIAs migrating to GB, as well as the limitations of the Tersoff and EAM IPs in describing these processes.
@@ 2.4 GB structure after defect migration
[…]
@@
This contrast in high-temperature robustness underscores the significance of training data. During the active learning approach, the MLIP-W accurately learned from ab initio data in these complex environments, capturing how defects are accommodated at the GB, thanks to a complete and rigorous description of the PES.
@@ 2.5 Descriptor-based analysis of defects
To further validate and analyze the performance of the MLIP-W in describing defect evolution at $\langle 110\rangle/W$$\langle 112\rangle$ GBs, we used the Fingerprinting and Visualization Analyzer of Defects (FaVAD) [40]. This approach is based on DVs and allows us to characterize defects based on their local atomic environments. Although the DeePMD-kit [43, 44, 69] generates its own DVs, their high dimensionality made them impractical for our 54,000-atom cell, and we instead use the multi-body descriptor framework called ‘Smooth Overlap of Atomic Positions’ (SOAP) [41] to compute a reduced set of descriptors (see Section 4.4). In this fingerprinting method, we measure the Euclidean distance, denoted as $d^{M}(T)$, between the DV of each atom in the defect sample, $\vec{q}^{\,i}$, and the mean DV of a defect-free (bulk) reference configuration at a certain temperature $T$, $\vec{v}(T)$. This distance difference between two local atomic environments forms the basis for a probabilistic interpretation of the results, expressed as $p\bigl{(}\vec{q}^{\,i}\mid\vec{v}(T)\bigr{)}$. We find that the distribution of these normalized distances is well approximated by a chi-squared distribution, $\chi^{2}_{K}$, with $K$ degrees of freedom. Mathematically, this is expressed as
@@
where $\Gamma$ is the Gamma function. In our analysis, as observed in Fig. S13, $K=4$ provides an optimal fit to the distributions [40]. This probabilistic framework allows defining a temperature-dependent threshold to differentiate genuine defects from thermal oscillations. To establish this threshold, first, the bulk reference DVs are computed at each simulation temperature. Since defects in BCC materials predominantly form $\langle 111\rangle$ dumbbell configurations [70], reference DVs for thermally equilibrated interstitials are also determined. Note that reference DVs are determined individually for each potential. By mapping distances relative to the bulk and interstitial reference DVs for a pure bulk configuration at a given temperature, a distribution that defines the bulk region in 2D space is obtained. Thus, when analyzing configurations containing defects or GBs, points outside this region are considered defects. This 2D methodology represents an advance over that used in Ref. [40], as it allows for a more accurate identification of crowdion-like interstitials, which are the most likely structures. Further details of this methodology, including how the area thresholds are determined, are given in Supplementary Note 5. In particular, Fig. S12 shows the distributions found for a 54,000-atom bulk system thermalized at each temperature and for each potential. The DV distributions reveal that while the MLIP-W broadens uniformly at higher temperatures, empirical IPs—particularly the Tersoff IP—develop bifurcated structures at 1000 K that are indicative of distinct structural motifs, a behavior likely attributable to an artifact of the potential. Exact thresholds can be found in Table S2 of the Supplementary Note 6.
@@
Once the bulk region is defined, atoms with DV values below the lower threshold are identified as dumbbell-like interstitials (see Fig. S12 for clarification). The remaining atoms—those near vacancy pairs, non-dumbbell-aligned interstitials, and atoms belonging to the GB—are grouped as unidentified defects. While a more detailed classification based on specific reference DV values for each defect type is feasible, such analysis is not needed for the goals of this study. On the other hand, identification of single vacancies and void formation was performed using Voronoi-based tessellation methods, as implemented in OVITO and described in previous works [71, 72]. This simple and robust method enables a quantitative estimation of local atomic volumes by dividing space into Voronoi polyhedra, thereby allowing detection of undercoordinated atoms and volumetric anomalies indicative of vacancy clusters or voids. Fig. 6 shows DV distributions (left) and simulation cells (right) for the high density of introduced defects case, both at 300 K and 1000 K. Atoms are color coded by environment: SIAs (light purple), vacancies (orange), and unidentified defects (dark purple), with bulk atoms omitted for the sake of clarity. Two new "sheets" emerge: the lower one for SIAs and the upper one for the GB and unidentified defects. In Fig. S13 of Supplementary Note 7, a detailed comparison of the DV distributions for bulk, pristine GB, and defected GB, each thermalized at 300 K, is presented. At 300 K, all potentials exhibit similar distributions, though the EAM IP shows a slight “V” shape. The MLIP-W produces fewer SIAs and defects, resulting in fewer points in the 3D cell. The EAM potential behaves similarly, whereas configurations produced with the Tersoff potential, as previously stated in Subsection 2.3, show substantial overlap between defect and thermally induced environments, making a clear defect identification difficult. At 1000 K, differences become more pronounced. The distribution found with the MLIP-W expands like in the bulk case, sharply reducing interstitials (by a factor of ten) and unidentified defects (by a factor of three) due to a broader bulk region filtering thermally vibrating atoms. The Tersoff IP shows an even steeper decline, from 4,129 defects at 300 K to only 185, although these numbers should be interpreted cautiously. EAM behaves similarly to the MLIP-W but introduces a third “sheet,” suggesting additional defective regions, i.e., atoms whose local environments deviate from the bulk and differ from the previous defect environments observed at lower temperatures. It also retains twice as many SIAs as the MLIP-W (253 vs. 138), likely due to fewer migration events.
@@ 2.6 Quantifying the temperature dependence of defect migration
Fig. 7 shows the average number of defects calculated with each IP as a function of temperature and initial defect density. Here, defects annihilated either by recombination with vacancies or migration into the GB region (treated as a defect sink) are not included in the final count. Full numerical values are provided in Supplementary Note 8, where Tables S3, S4, and S5 list the defect counts for the MLIP-W, the Tersoff, and the EAM potentials, respectively. At 300 K, in agreement with our previous observations, the MLIP-W consistently yields the lowest number of defects, whereas the Tersoff and the EAM potentials predict higher counts. Specifically, for the low density of introduced defects case, the number of defects calculated with the Tersoff and the EAM potentials is 1.7 and 1.2 times larger than that calculated with the MLIP-W, respectively. For the medium density of introduced defects case, the number of defects calculated with the Tersoff and the EAM IPs is 1.7 and 2.3 times higher than that calculated with the MLIP-W. Finally, for the high density of defects introduced case, the number of defects calculated with the Tersoff and the EAM IPs is 1.8 and 1.4 times larger than that calculated with the MLIP-W. These trends indicate that, at low temperatures, both the Tersoff and the EAM IPs yield more residual defects in the grain interior than the MLIP-W, evidencing the superior ability of the MLIP-W to capture defect migration and subsequent annihilation and/or reorganization at the GB. As temperature increases, the differences between the IPs become less pronounced. The EAM potential appears to follow a trend similar to the MLIP-W, albeit with consistently higher defect counts, whereas the slope of the Tersoff curves is considerably steeper, likely reflecting its inability to accurately capture defect behavior at high temperatures, resulting in defect counts even lower than those predicted by the MLIP-W.
@@
A natural way to validate the temperature dependence of these defect count curves is to interpret them in the light of the kinetic rate theory by Arrhenius [73, 62]. However, it is important to note that, as previously reported, deviations from Arrhenius behavior may occur at significantly higher temperatures than those considered in this study (i.e., temperatures ranging from approximately 60 to 75% of the melting temperature). At these high temperatures, the migration regime of interstitial defects transitions from discrete, infrequent hops to continuous Brownian-like motion driven by interactions with thermal phonons. This phenomenon takes place because the activation energy for SIA migration is so low (comparable to or smaller than thermal energy) that the defects do not remain localized for extended periods of time. Instead, they undergo continuous migration, resulting in a non-Arrhenius, linear temperature dependence of their diffusion coefficient at elevated temperatures [74, 75].
Nonetheless, in our study, the highest temperature explored (1000 K) is only about 27% of the melting temperature of tungsten (3695 K), falling well within the expected Arrhenius regime. For this reason and assuming that SIAs, once formed, diffuse to the GBs, that act as sinks for them, the annihilation rate or net decrease of SIAs at the GB with regard to the grain interior can be considered an effective measure of their migration. Such annihilation rate $v$ follows a classical Arrhenius form,
@@
where $\Delta n_{m}$ is the variation in the number of defects inside the grain (excluding vacancies) in a time $\Delta t$, $N$ is the total number of atoms, $S_{m}$ is the migration entropy, $E_{m}$ is the migration energy, and $k_{B}$ is the Boltzmann constant. The temperature-independent prefactor $\exp\bigl{(}S_{m}/k_{B}\bigr{)}$ represents the entropic contribution to the formation of an activated state, while $\exp\bigl{(}-E_{m}/k_{B}T\bigr{)}$ captures the energetic barrier to defect migration. Considering that $\Delta t$ is constant and equal to 100 ps in all simulations, we can fit Eq. (2) to our $\Delta n_{m}/N$ vs. $T$ data to extract $E_{m}$. We find that the MLIP-W yields a migration energy of $0.048\pm 0.001$ eV, which is in good agreement with the typical $\sim 0.05$ eV range reported for SIAs in W—for instance, 0.054 eV [62] or 0.046 eV [76]. For the EAM potential, the fits lead to a slightly lower migration energy value of $0.043\pm 0.001$ eV. The migration energy as fitted with the Tersoff potential has a much larger value than that calculated by the other potentials ($0.082\pm 0.008$ eV). These results are additional evidence of the inaccuracy of the Tersoff potential to distinguish defects from thermal noise. The good agreement of $E_{m}$ calculated with the MLIP-W and the reference data highlights its ability to capture defect kinetics and GB-mediated self-healing mechanisms in W under conditions where previously reported empirical IPs do not perform well.